Solve the system by substitution.
x + y = 12
y = 2x

Answer by checkley75(3666)

Y=2X NOW REPLACE Y IN THE OTHER EQUATION BY (2X) & SOLVE FOR X
X+2X=12
3X=12
X=12/3
X=4 ANSWER.
Y=2*4
Y=8 ANSWER.